# Efficiently Training 7B LLM with 1 Million Sequence Length on 8 GPUs

### v1

@@ Abstract.
Nowadays, Large Language Models (LLMs) have been trained using extended context lengths to foster more creative applications. However, long context training poses great challenges considering the constraint of GPU memory. It not only leads to substantial activation memory consumption during training, but also incurs considerable memory fragmentation. To facilitate long context training, existing frameworks have adopted strategies such as recomputation and various forms of parallelisms. Nevertheless, these techniques rely on redundant computation or extensive communication, resulting in low Model FLOPS Utilization (MFU). In this paper, we propose Memo, a novel LLM training framework designed for fine-grained activation memory management. Given the quadratic scaling of computation and linear scaling of memory with sequence lengths when using FlashAttention, we offload memory-consuming activations to CPU memory after each layer’s forward pass and fetch them during the backward pass. To maximize the swapping of activations without hindering computation, and to avoid exhausting limited CPU memory, we implement a token-wise activation recomputation and swapping mechanism. Furthermore, we tackle the memory fragmentation issue by employing a bi-level Mixed Integer Programming (MIP) approach, optimizing the reuse of memory across transformer layers. Empirical results demonstrate that Memo achieves an average of $2.42\times$ and $2.26\times$ MFU compared to Megatron-LM and DeepSpeed, respectively. This improvement is attributed to Memo’s ability to minimize memory fragmentation, reduce recomputation and intensive communication, and circumvent the delays associated with the memory reorganization process due to fragmentation. By leveraging fine-grained activation memory management, Memo facilitates efficient training of 7B LLM with 1 million sequence length on just 8 A800 GPUs, achieving an MFU of 52.30%.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Since the advent of ChatGPT (OpenAI, 2023), Large Language Models (LLMs) have demonstrated remarkable proficiency in comprehending and generating natural language texts. Besides revolutionizing the field of language processing, which encompasses translation (Zhu et al., 2023), coding (Rozière et al., 2023; Guo et al., 2024b), etc., transformer-based LLMs have also found applications in multi-modal scenarios, such as image processing (Dosovitskiy et al., 2021; Peebles and Xie, 2023), video stream analysis (Ruan and Jin, 2022), and AI for science (Bi et al., 2023; DBL, 2023). To accommodate novel applications that require lengthy contexts, LLMs have developed to support long context input, from 2K-4K (Touvron et al., 2023; Taori et al., 2023) to 32K (Jiang et al., 2023; Together.ai, 2023), 128K (OpenAI, 2023; Fu et al., 2024), or even millions of tokens (AI, 2024; Cloud, 2024; Liu et al., 2024). Considering the extrapolation problem (Press et al., 2022; Liu et al., 2023a), which refers to the decline in LLM performance when input sequences exceed the training length, it is necessary to conduct long context training (Jacobs et al., 2023; Chen et al., 2024; Fang and Zhao, 2024) or fine-tuning (Peng et al., 2023; Ding et al., 2024) to facilitate long sequence inference. Beyond natural language processing, increasing the context length is also essential across diverse domains, including video processing (Zheng et al., 2024), protein properties prediction (Chandra et al., 2023), weather forecasting (Nguyen et al., 2023), and health care (Li et al., 2022).
@@
However, training LLMs with long sequence lengths poses a significant challenge for GPU memory. During training, a large amount of activations (i.e., the intermediate results computed in the forward pass) must be stored for gradient computation during the backward pass, resulting in substantial memory consumption. Typically, it is well known that the self-attention module in the transformer architecture has a quadratic computation and memory complexity w.r.t. the sequence length. FlashAttention (Dao et al., 2022; Dao, 2023), now a standard technique for attention computation in LLM training, accelerates computation and shrinks the memory complexity to be linear w.r.t. the sequence length by scheduling memory I/O and recomputing necessary components during the backward pass. Except for attention, the remaining activation memory also scales proportionally with the sequence length, which can become quite large in long context scenarios. For instance, training a GPT model with 7B parameters on a sequence length of 1 million can lead to an activation memory of 4096GB, far exceeding the memory capacity of commonly used accelerators (e.g. 80GB for an NVIDIA H100/A100 GPU).
@@
Moreover, the memory fragmentation issue makes this situation even worse. Besides storing the skeletal activations for the backward pass, there are also tremendous transient activations that are temporarily generated during computation (we will formally categorize the two kinds of activations in Section 3). Such transient activations have distinct data life cycles and usually lead to frequent allocation and deallocation of GPU memory. Currently, most LLM training systems are built on top of PyTorch (Paszke et al., 2019), including Megatron-LM (Shoeybi et al., 2019) and DeepSpeed (Rasley et al., 2020). PyTorch employs a caching memory allocator designed to reduce the costly “cudaMalloc” and “cudaFree” operations by caching and reusing allocated memory blocks. However, the frequent memory (de)allocation requests in the caching allocator result in significant memory fragmentation (Guo et al., 2024a). This issue becomes more severe in long context training, considering the fact that the (de)allocated memory blocks are significantly larger than those in normal tasks. Memory fragmentation not only leads to Out-Of-Memory (OOM) error but also significantly hinders training efficiency because of the frequent triggering of the PyTorch memory reorganization process, which involves calls to expensive “cudaFree” and “cudaMalloc” to release cached blocks and reclaim GPU memory. Figure 1(a) illustrates an example of GPU memory fragmentation. At the peaks of the curves, there is more than 4GB memory reserved but not allocated. However, when the training task tries to allocates 4GB memory, the allocator fails to find a continuous memory space to fulfill the allocation request. Consequently, it necessitates to invoke a series of “cudaFree” and “cudaMalloc” to reorganize memory, which blocks GPU computation.
(a) GPU memory consumption.
(b) One-layer forward time consumption.
@@
Observation 1: Opportunity for activation swapping. In the field of deep learning training, to reduce the peak memory consumption caused by skeletal activations, activation recomputation  (Chen et al., 2016; Kirisame et al., 2021; Korthikanti et al., 2022) and swapping (Rhu et al., 2016; NVIDIA, 2024b) are two well-known memory reduction techniques that trade time to save memory. [^1] Parallelism techniques like sequence parallelism (Korthikanti et al., 2022; Jacobs et al., 2023) and context parallelism (Liu et al., 2023b; Li et al., 2023; NVIDIA, 2024a) are also compelling approaches to reduce memory at the price of extra communication overhead. Our work is compatible with these parallelism techniques. Typically, both of them reduce memory consumption at the price of extra time cost. For one thing, the activation recomputation technique discards the skeletal activations in the forward pass and later recomputes them in the backward pass, leading to extra computation cost. For another, the swapping technique offloads the activations to CPU memory in the forward pass to relieve the GPU memory pressure, and later fetches them back to GPU memory in the backward pass, incurring the overhead of data transmission between CPU and GPU memory. Contemporary, mainstream LLM training frameworks such as Megatron-LM and DeepSpeed prefer activation recomputation to swapping, which is due to the fact that the GPU computing ability has a far more rapid growth than the connectivity between CPU and GPU memory in the past few years (see Section 2.2 for details).
@@
However, we find that the situation is a bit different in long context training of LLMs. Denote $s$ as the sequence length. The computation complexity of one transformer layer is $O(s^{2})$, while the activation memory complexity is $O(s)$ thanks to FlashAttention. During GPU computation, we can leverage the idle CPU-GPU bandwidth, offloading activations to CPU memory during the forward pass, and fetching the activations during the backward pass. As the sequence length increases, there is greater potential for overlapping computation and communication, given that their time requirements scale quadratically and linearly with the sequence length, respectively. As shown in Figure 1(b), eventually, after reaching a specific sequence length (192K in this case), the transmission of activations can be fully overlapped with GPU computation.
However, in practice, there is limited chance to completely swap all activations. On the one hand, extremely long training data is rare, and most of the time we need to train on data that doesn’t fully overlap the activation transmission and the computation. On the other hand, offloading all activations may cause CPU OOM issues — the CPU memory is responsible for storing all activations from all GPUs on the same machine, but the current CPU memory is typically several terabytes, which is insufficient for very long sequence lengths. Considering the above challenges, we introduce a fine-grained activation recomputation and swapping mechanism to manage the skeletal activations. We consider both tensor-level and token-level activation management. For each layer, following previous works (NVIDIA, 2024b; Li et al., 2023), we consistently offload two kinds of activation tensors, the input of each transformer layer and the output of FlashAttention, to CPU memory. For other activation tensors, we only offload a fraction (denoted as $\alpha$) of tokens, and recompute the rest part during the backward pass. We model the time cost of activation recomputation and transmission and determine the fraction $\alpha$ through a well-formulated linear programmin problem, which aims to maximize offloading activations without impeding GPU computation or causing CPU OOM issues. During the backward pass, prefetching activations can also overlap with GPU computation, because the backward computation is typically twice as much as the forward computation. With both tensor-level and token-level activation management, we make full use of the idle bandwidth and minimize the recomputation overhead to improve the overall efficiency.
@@
Figure 2. An overview of Memo. We devise a fine-grained recomputation and swapping mechanism to manage the skeletal activations for backward propagation, and leverage a bi-level memory planning to reuse the memory space of transient activations across the transformer layers.
@@
Observation 2: Deterministic memory (de)allocation pattern across iterations and layers. The memory fragmentation mainly comes from frequent and irregular memory (de)allocation requests. However, we observe that, typical LLM training adheres to a deterministic computation process across iterations and layers. All transformer layers in an LLM are identical, and each training iteration involves the same computation. While the general-purpose caching allocator is designed for dynamic computation routines, training LLMs can be conceptualized as static computation graphs (Ansel et al., 2024), which have identical structures across layers. This provides an opportunity to design static planning for each layer and reuse the allocated memory of each layer, thereby mitigating memory fragmentation.
@@
To enhance memory utilization while minimizing fragmentation, we leverage a hierarchical Mixed Integer Programming (MIP) technique to tackle the memory planning problem. Before training, we profile the memory (de)allocation requests of one training iteration, then use MIP to solve an optimized memory plan for a single transformer layer. Since each layer’s memory request are identical, the entire memory block for one layer can be directly reused for the subsequent identical layer. Considering each transformer layer’s memory block as a single memory allocation request, we further solve another MIP problem that plans memory allocation for the entire LLM training, including the initial embedding layer, all transformer layers, and the final classifier layer. We only need to solve the problem once before the actual training, since all iterations can utilize the same memory plan. The near-optimal memory plan eliminates the fragmentation issue and avoids PyTorch’s time-consuming memory reorganization mechanism.
@@
Putting them together, in response to the activation memory challenge in long context training, we propose Memo, an LLM training framework with fine-grained tensor memory management. We consider the challenge as an activation data management problem. To make full use of the idle CPU-GPU bandwidth during training with different sequence lengths, we introduce a token-wise fine-grained activation recomputation and swapping strategy. We employ a bi-level hierarchical MIP technique to solve the memory planning problem and eliminate memory fragmentation. To the best of our knowledge, this is the first training framework that enables efficient training of a 7B LLM on 8 GPUs with a sequence length of 1 million.
@@
  • We employ a bi-level hierarchical MIP technique to solve the memory planning problem and significantly mitigate memory fragmentation.
  • We evaluate Memo through extensive experiments, and demonstrate an average of $2.42\times$ and $2.26\times$ MFU compared to Megatron-LM and DeepSpeed, respectively. Additionally, Memo is the first framework that enables the efficient training of 7B LLM with 1 million context length on only 8 A800 GPUs.

### v2

# Memo: Fine-grained Tensor Management For Ultra-long Context LLM Training

@@ Abstract.
Nowadays, Large Language Models (LLMs) have been trained using extended context lengths to foster more creative applications. However, long context training poses great challenges considering the constraint of GPU memory. It not only leads to substantial activation memory consumption during training, but also incurs considerable memory fragmentation. To facilitate long context training, existing frameworks have adopted strategies such as recomputation and various forms of parallelisms. Nevertheless, these techniques rely on redundant computation or extensive communication, resulting in low Model FLOPS Utilization (MFU). In this paper, we propose Memo, a novel LLM training framework designed for fine-grained activation memory management. Given the quadratic scaling of computation and linear scaling of memory with sequence lengths when using FlashAttention, we offload memory-consuming activations to CPU memory after each layer’s forward pass and fetch them during the backward pass. To maximize the swapping of activations without hindering computation, and to avoid exhausting limited CPU memory, we implement a token-wise activation recomputation and swapping mechanism. Furthermore, we tackle the memory fragmentation issue by employing a bi-level Mixed Integer Programming (MIP) approach, optimizing memory reuse across transformer layers. Empirical results demonstrate that Memo achieves an average of $1.97\times$ and $1.80\times$ MFU compared to Megatron-LM and DeepSpeed, respectively. This improvement is attributed to Memo’s ability to minimize memory fragmentation, reduce recomputation and intensive communication, and circumvent the delays associated with the memory reorganization process due to fragmentation. By leveraging fine-grained activation memory management, Memo facilitates efficient training of 7B LLM with 1 million sequence length on just 8 A800 GPUs, achieving an MFU of 52.30%.
Large Language Model, Long Context Training, Tensor Management, Data Rematerialization
copyright: acmlicensedjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 1 (SIGMOD)article: 53publicationmonth: 2doi: 10.1145/3709703isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Distributed computing methodologies
@@ 1. Introduction
Since the advent of ChatGPT (OpenAI, 2023), Large Language Models (LLMs) have demonstrated remarkable proficiency in comprehending and generating natural language texts. Besides revolutionizing the field of language processing, which encompasses translation (Zhu et al., 2024), coding (Rozière et al., 2023; Guo et al., 2024b; Zhang et al., 2024b), etc., transformer-based LLMs have also found applications in multi-modal scenarios, such as image processing (Dosovitskiy et al., 2021; Peebles and Xie, 2023), video stream analysis (Ruan and Jin, 2022), and AI for science (Bi et al., 2023; AI4Science and Quantum, 2023). To accommodate novel applications that require lengthy contexts (Zhao et al., 2024), LLMs have developed to support long context input, from 2K-4K (Touvron et al., 2023; Taori et al., 2023) to 32K (Jiang et al., 2023; Together.ai, 2023), 128K (OpenAI, 2023; Fu et al., 2025), or even millions of tokens (AI, 2024; Cloud, 2024; Liu et al., 2024b). Considering the extrapolation problem (Press et al., 2022; Liu et al., 2024a), which refers to the decline in LLM performance when input sequences exceed the training length, it is necessary to conduct long context training (Jacobs et al., 2023; Chen et al., 2024; Fang and Zhao, 2024) or fine-tuning (Peng et al., 2024; Ding et al., 2024) to facilitate long sequence inference. Beyond natural language processing, increasing the context length is also essential across diverse domains, including video processing (Zheng et al., 2024), protein properties prediction (Chandra et al., 2023), weather forecasting (Nguyen et al., 2023), and health care (Li et al., 2022).
@@
Maximizing system performance with limited memory is a common and significant challenge in the data management community. Within this context, training LLMs with long sequence lengths poses difficulties due to restricted GPU memory. During training, a large amount of activations [^1] In neural network training, the outputs of operators are referred to as activations. Some of these, termed “skeletal activations” in this work, must be stored for gradient computation during the backward pass, while others are termed “transient activations”. More details are provided in Section 3. must be stored for gradient computation during the backward pass, resulting in substantial memory consumption. Typically, it is well known that the self-attention module in the transformer architecture has a quadratic computation and memory complexity w.r.t. the sequence length. FlashAttention (Dao et al., 2022; Dao, 2024), now a standard technique for attention computation in LLM training, accelerates computation and shrinks the memory complexity to be linear w.r.t. the sequence length by scheduling memory I/O and recomputing necessary components during the backward pass. Except for attention, the remaining activation memory also scales linearly with the sequence length, which can become quite large in long context scenarios. For instance, training a GPT model with 7B parameters on a sequence length of 1 million can lead to an activation memory of 4096GB, far exceeding the memory capacity of commonly used accelerators (e.g. 80GB for an NVIDIA H100/A100 GPU).
@@
Moreover, dynamic memory allocators inherently face the issue of memory fragmentation due to frequent allocation and deallocation, which complicates efficient data management. Besides storing the skeletal activations for the backward pass, there are also tremendous transient activations that are temporarily generated during computation (we will formally categorize the two kinds of activations in Section 3). Such transient activations have distinct data life cycles and usually lead to frequent allocation and deallocation of GPU memory. Currently, most LLM training systems are built on top of PyTorch (Paszke et al., 2019), including Megatron-LM (Shoeybi et al., 2019) and DeepSpeed (Rasley et al., 2020). PyTorch employs a caching memory allocator designed to reduce the costly “cudaMalloc” and “cudaFree” operations by caching and reusing allocated memory blocks. However, the frequent memory (de)allocation requests in the caching allocator result in significant memory fragmentation (Guo et al., 2024a). This issue becomes more severe in long context training, considering the fact that the (de)allocated memory blocks are significantly larger than those in normal tasks. Memory fragmentation not only leads to Out-Of-Memory (OOM) error but also significantly hinders training efficiency because of the frequent triggering of the PyTorch memory reorganization process, which involves calls to “cudaFree” and “cudaMalloc” to release cached blocks and reclaim GPU memory. Figure 1 illustrates an example of GPU memory fragmentation. At the peaks of the curves, there is more than 4GB memory reserved but not allocated. However, when the training task tries to allocates 4GB memory, the allocator fails to find a continuous memory space to fulfill the allocation request. Consequently, it necessitates invoking a series of “cudaFree” and “cudaMalloc” to reorganize memory, which blocks GPU computation.
@@
In the data management domain, when high-bandwidth memory is limited, rematerialization is a typical technique to free up memory by releasing data that is not immediately needed and reconstructing data structures on demand. For instance, Spark (Zaharia et al., 2016) supports rematerializing Resilient Distributed Datasets (RDDs) (Zaharia et al., 2012) through recomputation or swapping [^2] Swapping refers to moving data between different levels of memory hierarchy. More details are provided in Section $2.$ between CPU memory and disk storage when RAM is insufficient. Faiss (Johnson et al., 2021), a widely used vector database, utilizes a CPU-GPU memory hierarchy to accommodate the high memory requirement. Additionally, the dynamic view materialization solution (Phan and Li, 2008) employs an LRU cache to manage materialized views, and rematerializes cache-missed views upon needed. These methods leverage recomputation and swapping strategies to efficiently manage data structures within the memory hierarchy. Similarly, in deep learning training, to reduce the peak memory consumption caused by skeletal activations, activation recomputation  (Chen et al., 2016; Kirisame et al., 2021; Korthikanti et al., 2022) and swapping (Rhu et al., 2016; NVIDIA, 2024b) are also widely adopted.  [^3] Parallelism techniques like sequence parallelism (Korthikanti et al., 2022; Jacobs et al., 2023) and context parallelism (Liu et al., 2023; Li et al., 2023a; NVIDIA, 2024a) are also compelling approaches to reduce memory at the price of extra communication overhead. Our work is compatible with these parallelism techniques. Typically, both of them reduce memory consumption at the price of extra time cost. The activation recomputation technique discards some activations in the forward pass and later recomputes them in the backward pass, leading to extra computation cost. The swapping technique offloads the activations to CPU memory in the forward pass to relieve the GPU memory pressure, and later fetches them back to GPU memory in the backward pass, incurring the overhead of data transmission between CPU and GPU memory.
@@
Observation 1: Opportunity for activation swapping. Contemporary mainstream LLM training frameworks such as Megatron-LM and DeepSpeed prefer activation recomputation to swapping, which is due to the fact that the GPU computing ability has a far more rapid growth than the connectivity between CPU and GPU memory in the past few years (see Section 2.2 for details). However, we find that the situation is a bit different in long context training of LLMs. Denote $s$ as the sequence length. The computation complexity of one transformer layer is $O(s^{2})$, while the activation memory complexity is $O(s)$ thanks to FlashAttention. During GPU computation, we can leverage the idle CPU-GPU bandwidth, offloading activations to CPU memory during the forward pass, and fetching the activations during the backward pass. As the sequence length increases, there is greater potential for overlapping computation and communication, given that their time requirements scale quadratically and linearly with the sequence length, respectively. As shown in Figure 1, eventually, after reaching a specific sequence length (192K in this case), the transmission of activations can be fully overlapped with GPU computation.
@@
Figure 2. An overview of Memo. We devise a fine-grained recomputation and swapping mechanism to manage the skeletal activations for backward propagation, and leverage a bi-level memory planning method to reuse the memory space of transient activations across the transformer layers.
@@
However, in practice, there is limited chance to swap all activations. On the one hand, extremely long training data is rare, and most of the time we need to train on data that doesn’t fully overlap the activation transmission and the computation. On the other hand, offloading all activations may cause CPU OOM issues — the CPU memory is responsible for storing all activations from all GPUs on the same machine, but the current CPU memory is typically several terabytes, which is insufficient for very long sequence lengths. Considering the above challenges, we introduce a fine-grained activation recomputation and swapping mechanism to manage the skeletal activations. We consider both tensor-level and token-level activation management. For each layer, following previous works (NVIDIA, 2024b; Li et al., 2023a), we consistently offload two activation tensors, the input of each transformer layer and the output of FlashAttention, to CPU memory. For other activation tensors, we only offload a fraction (denoted as $\alpha$) of tokens, and recompute the rest part during the backward pass. We model the time cost of activation recomputation and transmission and determine the fraction $\alpha$ through a well-formulated linear programming problem, which aims to maximize offloading activations without impeding GPU computation or causing CPU OOM issues. During the backward pass, prefetching activations can also overlap with GPU computation, because the backward computation is typically twice as much as the forward computation. With both tensor-level and token-level activation management, we make full use of the idle bandwidth and minimize the recomputation overhead to improve the overall efficiency.
Regarding memory fragmentation, research has utilized the characteristics of targeted workloads to analyze and resolve the issue, such as experimentally analyzing the impact of memory allocation on high-performance query engines (Durner et al., 2019) and addressing persistent memory fragmentation with efficient defragmentation algorithms (Oukid et al., 2017). In the same vein, for long-sequence LLM training, we aim to leverage the specific characteristics of LLM training to address memory fragmentation.
@@
Observation 2: Deterministic memory (de)allocation pattern across iterations and layers. In long sequence LLM training, the memory fragmentation mainly comes from frequent and irregular memory (de)allocation requests. However, we observe that, typical LLM training adheres to a deterministic computation process across iterations and layers. All transformer layers in an LLM are identical, and each training iteration involves the same computation. While the general-purpose caching allocator is designed for dynamic computation routines, training LLMs can be conceptualized as static computation graphs (Ansel et al., 2024), which have identical structures across layers. This provides an opportunity to design static planning for each layer and reuse the allocated memory of each layer, thereby mitigating memory fragmentation.
@@
To enhance memory utilization while minimizing fragmentation, we leverage a hierarchical Mixed Integer Programming (MIP) technique to tackle the memory planning problem. Before training, we profile the memory (de)allocation requests of one training iteration, then use MIP to solve an optimized memory plan for a single transformer layer. Since the memory requests of the transformer layers are identical, the entire memory block for one layer can be directly reused for the subsequent identical layer. Considering each transformer layer’s memory block as a single memory allocation request, we further solve another MIP problem that plans memory allocation for the entire LLM training, including the initial embedding layer, all transformer layers, and the final classifier layer. We only need to solve the problem once before the actual training, since all iterations can utilize the same memory plan. The near-optimal memory plan eliminates the fragmentation issue and avoids PyTorch’s time-consuming memory reorganization mechanism.
@@
Putting them together, in response to the activation memory challenge in long context training, we propose Memo, an LLM training framework with fine-grained tensor memory management. We consider the challenge as an activation data management problem, and draw inspiration from the data rematerialization and memory defragmentation techniques to address the challenge. Figure 2 presents an overview of Memo. To make full use of the idle CPU-GPU bandwidth during training with different sequence lengths, we introduce a token-wise fine-grained activation recomputation and swapping strategy. We employ a bi-level hierarchical MIP technique to solve the memory planning problem and eliminate memory fragmentation. To the best of our knowledge, this is the first training framework that enables efficient training of a 7B LLM on 8 GPUs with a sequence length of 1 million.
@@
  • We employ a bi-level MIP technique to solve the memory planning problem and significantly mitigate memory fragmentation.
  • We evaluate Memo through extensive experiments, and demonstrate an average of $1.97\times$ and $1.80\times$ improvement in terms of MFU [^4] MFU (Model FLOPs Utilization) is a widely used efficiency metric that evaluates how well the accelerators are utilized in model training. It is calculated as the ratio of the observed throughput to the theoretical throughput which assumes the hardware operates at peak FLoating-point Operations Per Second (FLOPS) (Pope et al., 2023) . More details are provided in Section 5.1. compared to Megatron-LM and DeepSpeed, respectively. Additionally, Memo is the first framework that enables the efficient training of 7B LLM with 1 million context length on only 8 A800 GPUs.